# Generalizing Visual Question Answering from Synthetic to Human-Written Questions via a Chain of QA with a Large Language Model

### v1

@@ 5.1 Benchmarks on CLEVR-Human
Table 1 compares accuracy between the general VLMs, template-based, and finetuned VQA models. Overall, the general VLMs reported lower accuracy than the template-based VQA models, except for the latest GPT-4-vision (e.g., 60.1% for GPT-4-vision vs. 59.9% for MDETR). CoQAH achieved the highest accuracy with large gaps compared to all the general VLMs and template-based VQA models (e.g., 74.3% accuracy for CoQAH vs. 59.9% for MDETR). Although there are still some performance gaps between CoQAH and the finetuned models (e.g., 74.3% for CoQAH vs. 81.7% for finetuned MDTER), it is very difficult to obtain a large number of human-written QA pairs in general to finetune the model accordingly.
@@ 5.2 Benchmarks on VQA-RAD and SLAKE
[…]
@@ 5.3 Interpreting Reasons for Answers
Figure 4 shows some examples that we asked an LLM in CoQAH to interpret the reasons for answers. As shown in the examples, CoQAH successfully delineated the rationale behind the answers based on the information collected through the dialogues. For example, for the question of whether there are two cyan balls of the same size (Figure 4a), CoQAH reasonably explained the reason (one is large and one is small) for the answer (No) after checking the size of the cyan balls (Is there a large cyan ball? and Is there a small cyan ball?). More examples can be found in Supplementary Figure 8 (for 3D-rendered images) and 9 (for chest X-rays).
@@ 5.4 Effect of the Maximum Number of Questions
Table 3 shows the accuracy of CoQAH on the CLEVR-Human when we changed the maximum number of questions to be queried. The maximum accuracy was achieved when the maximum number of questions was 20, and even if we increased the number to 30, the accuracy was the same.
@@
Table 3: Accuracy on CLEVR-Human validation dataset according to the changes of the maximum number of questions to be queried in CoQAH. We only used 5% of the validation data for the ablation study.
5.5 Ablation Study
The results of the ablation study are summarized in Table 4. When we ablated EUH in CoQAH, the accuracy on CLEVR-Human was slightly decreased (75.8% for CoQAH vs. 74.7% without EUH). When ablating the step-by-step QA process in CoQAH, the accuracy was degraded mainly (75.8% for CoQAH vs. 63.1% without the process), indicating the importance of this process to derive the answers correctly.Finally, when we ablated the template format description in the task instruction, the accuracy was dramatically degraded (75.8% for CoQAH vs. 35.8% without the format description). This means that the questions from the LLM must conform to the template format so the VQA model can understand those questions.
@@ 5.5 Ablation Study
Table 4: Comparison of accuracy on CLEVR-Human by ablating either EUH or step-by-step QA process. Ablating the QA process degraded the accuracy significantly. Ablating the template format description dramatically degraded the accuracy, meaning that the questions from the LLM must conform to the template format.
@@ 5.6 Error Analysis of CoQAH
The most significant portion of errors (48%) occurred when the LLM failed to ask proper questions to derive correct answers (e.g., Supplementary Figure. 8e). The second largest portion (25%) was due to incorrect responses from the VQA model during the QA process (e.g., Supplementary Figure. 8d). Some errors (11%) dedicated to the questions where the LLM was impossible to answer by querying template-based questions only (e.g., Supplementary Figure. 8f). Other errors (15%) included failures of following instructions, data format (e.g., The answer should be in {}), and so on.
@@ 6 Conclusion
In this study, we proposed a novel CoQAH methodology that answers human-written questions via a reasoning process based on a dialogue. Throughout the extensive experiments, we have shown that CoQAH achieved state-of-the-art performance on various datasets of different images (3D-rendered and X-ray) without finetuning using the human-written questions. We believe that this new method potentially promotes the wide applications of AI for visual question answering, improving the practicality of AI in real environment.
@@ 7 Limitation
First, we could only test a limited number of general VLM models (LLaVA Liu et al. (2023) and GPT-4-Vision OpenAI (2023)) for CLEVR-Human since, currently, many LLMs do not support taking an image as an input. Second, we reported the accuracy of the AI models we ran on CLEVR-Human validation data because the test set was unavailable. Third, even though we tried our best to optimize the task instructions for the general VLM and medical foundation models, those might not be the best due to the lack of methods to tune the instructions. Fourth, we only validated CoQAH using two types of images (3D-rendered and chest X-ray) because of a limited number of template-based datasets.
Ethical Statement
There are no ethical issues.
Acknowledgments
Acknowledgments
@@
FiLM Perez et al. [2018] 97.7
MAC Hudson and Manning [2018] 98.9
MDETR Kamath et al. [2021] 99.7
Table 1: The performance of VQA models on the CLEVR testing dataset. Regardless of the models, the accuracy values were very high ($>$ 97%), meaning that those VQA models can precisely answer the template-based questions. We used the MDETR model as a template-based VQA model for CoQAH.
@@
EKIDG Hu et al. [2023] 0.624 0.541 0.477 0.422 0.337 0.645 1.893
OFA-MIMIC Wang et al. [2022] 0.662 0.588 0.528 0.473 0.362 0.716 2.332
[…]
Figure 1: Template format for the chest X-ray VQA dataset, MIMIC-Diff-VQA. A question is only composed of two parts, $\langle type\rangle$ and $\langle abnormality\rangle$. Similar to the case of the CLEVR dataset, we asked an LLM to select an option available for each $\langle type\rangle$ and $\langle abnormality\rangle$ to generate a template-based question.
Figure 2: Task instruction of CoQAH for CLEVR.
Figure 3: Task instruction of CoQAH for MIMIC-Diff-VQA.
Figure 4: Task instruction of LLaVA for CLEVR-Human. (a) In the first stage, we asked LLaVA to produce an answer as a single word or phrase. (b) In some cases, LLaVA provided answers of the same meaning but different expressions. Therefore, in the second stage, we corrected them by asking LLaVA to choose one of the candidates (e.g., round for the first answer vs. sphere for the corrected answer). Note that when we gave all options for the answer at once, the performance of LLaVA was primarily degraded.
Figure 5: Task instruction of GPT-4-Vision for CLEVR-Human. We provided a set of answers.
Figure 6: Task instruction of Med-Flamingo for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
Figure 7: Task instruction of MedVInT-TD for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
@@
Table 3: Table summarizing the answers of AI we substituted when calculating accuracy. We changed them because those words had the same meaning but different expressions, resulting in misleading performance metrics.
Figure 8: Example dialogues of CoQAH for CLEVR-Human. (a)-(c) The cases show that CoQAH correctly answered the questions and explained the reasons. (d) The answer was wrong because the template-based VQA model failed to give correct information. (e)-(f) CoQAH failed since the large language model could not ask relevant questions to derive correct answers.
Figure 9: Example dialogues of CoQAH for VQA-RAD and SLAKE. (a)-(c) CoQAH successfully answered the given user questions. (d) The VQA model failed to give the correct answer, and CoQAH failed accordingly. (e) The case shows that the large language model logically derived the answer even though the answer was incorrect. (f) Since the VQA model could not detect the lesion, CoQAH also gave the wrong answer.

### v3

@@ 5.1 Benchmarks on CLEVR-Human
Table 1 compares accuracy between the general VLMs, template-based, and finetuned VQA models. Overall, the general VLMs reported lower accuracy than the template-based VQA models, except for the latest GPT-4-vision (e.g., 60.1% for GPT-4-vision vs. 59.9% for MDETR). CoQAH achieved the highest accuracy with large gaps compared to all the general VLMs and template-based VQA models (e.g., 74.3% accuracy for CoQAH vs. 59.9% for MDETR). There are still some performance gaps between CoQAH and the finetuned models (e.g., 74.3% for CoQAH vs. 81.7% for finetuned MDTER). However, it is worth noting that obtaining a large number of human-written QA pairs and finetuning a model is very difficult for most VQA tasks in general, and our framework substantially improves VQA performance using only synthetic QA pairs that are much easier to collect.
@@ 5.2 Benchmarks on VQA-RAD and SLAKE
Table 2 summarizes the performance of the medical foundation and template-based VQA models for VQA-RAD and SLAKE. Among all the models, CoQAH reported the highest accuracy in the closed-form questions (e.g., VQA-RAD: 67.5% for CoQAH vs. 59.5% for OFA-MIMIC, SLAKE: 73.9% for CoQAH vs. 69.4% for OFA-MIMIC), and also the highest LAVEGPT-4 in the open-form questions (e.g., VQA-RAD: 0.302 for CoQAH vs. 0.274 for MedVInT-TD, SLAKE: 0.425 for CoQAH vs. 0.396 for MedVInT-TD). There were huge gaps between the LAVEGPT-4 of the OFA-MIMIC and CoQAH, meaning that the chain of QA process in CoQAH indeed improved the correctness of the answers significantly.
@@ 5.3 Interpreting Reasons for Answers
Since the CoQAH method utilizes an LLM to understand the context and finally answer the user question, we can ask the rationale of the final answer to the LLM (i.e., interpreting the AI model’s behavior), simply querying ‘why?’. For a few complex questions, we asked the LLM to explain the reasons to check whether it had logically driven their conclusions.
As shown in the examples of Figure 4, CoQAH successfully delineated the rationale behind the answers based on the information collected through the dialogues. For example, for the question of whether there are two cyan balls of the same size (Figure 4a), CoQAH reasonably explained the reason (one is large and one is small) for the answer (No) after checking the size of the cyan balls (Is there a large cyan ball? and Is there a small cyan ball?). More examples can be found in Supplementary Figure 8 (for 3D-rendered images) and 9 (for chest X-rays).
@@ 5.4 Effect of the Maximum Number of Questions
The maximum number of questions to be queried is important parameter since it can directly affect the quality of the answers from CoQAH. Therefore, we measured the accuracy of CoQAH on the CLEVR-Human validation dataset (5% used), changing that number from 0 to 30. As shown in Table 3, The maximum accuracy was achieved when the maximum number of questions was 20, and even if we increased the number to 30, the accuracy was the same.
5.5 Ablation Study
We conducted an ablation study by excluding several components in CoQAH and measured the accuracy based on the CLEVR-Human validation data (5% used). The results of the ablation study are summarized in Table 4. When we ablated EUH in CoQAH, the accuracy on CLEVR-Human was slightly decreased (75.8% for CoQAH vs. 74.7% without EUH). When ablating the step-by-step QA process in CoQAH by enforcing the LLM to ask all the questions simultaneously (ask me up to 20 questions all at once), the accuracy was degraded mainly (75.8% for CoQAH vs. 63.1% without the process), indicating the importance of this process to derive the answers correctly. Finally, when we ablated the format description in the task instruction by letting LLM generate any form of questions during the chain of QA process, the accuracy was dramatically degraded (75.8% for CoQAH vs. 35.8% without the format description). This means that the questions from the LLM must conform to the template format so the VQA model can understand those questions. Additionally, to investigate the effect of few-shot prompting on CoQAH, we compared the performance of CoQAH in zero-shot, one-shot, and two-shot settings. Dialogues between the VQA model and the LLM were provided as few-shot exemplars to the LLM. The results are shown in Table  5. The performance increased when few-shot exemplars were given, but increasing the number of exemplars from one to two did not further improve performance.
Table 3: Accuracy on CLEVR-Human validation dataset according to the changes of the maximum number of questions to be queried in CoQAH. We only used 5% of the validation data for the ablation study. Until the maximum number of questions reached 20, accuracy increased as the number of questions increased. However, the accuracy remained the same when the maximum number of questions was set at 20 and 30.
@@ 5.5 Ablation Study
Table 4: Comparison of accuracy on CLEVR-Human by ablating the step-by-step QA process, EUH, or the template format description. Ablating the QA process degraded the accuracy significantly. Ablating the format description dramatically degraded the accuracy, meaning that the questions from the LLM must conform to the template format.
@@
Table 5: The performance of CoQAH in zero-shot, one-shot, and two-shot settings. Dialogues between the VQA model and the LLM were provided as few-shot exemplars to the LLM. Performance increased when few-shot exemplars were provided, but adding more exemplars, from one to two, did not result in any further improvement.
Prompt Setting Zero-shot One-shot Two-shot
Acc. on 75.8 80.0 79.4
CLEVR-Human (%)
@@ 5.6 Error Analysis of CoQAH
We performed an error analysis by investigating the dialogues where CoQAH failed to answer correctly (5% CLEVR-Human validation used) and categorized them according to the reasons (e.g., incorrect answers from VQA model). The results of the error analysis are shown in Table 6.
The most significant portion of errors (48%) occurred when the LLM failed to query key questions or made a hasty conclusion after asking some irrelevant questions (e.g., Supplementary Figure 8e). It means that the LLM still has shown a limited reasoning ability during the QA process for some complex questions. The second largest portion (25%) was due to incorrect responses from the VQA model during the QA process (e.g., Supplementary Figure 8d). Some errors (11%) dedicated to the questions where the LLM was impossible to answer by querying template-based questions only (e.g., Supplementary Figure 8f). Other errors (15%) included failures of following instructions, data format (e.g., The answer should be in {}), and so on.
Table 6: Results of the error analysis for CoQAH. We investigated the dialogues where CoQAH failed to answer correctly and categorized them according to the reasons. The most significant portion of errors occurred when the LLM failed to query key questions or made a hasty conclusion after asking some irrelevant questions, showing a limited reasoning ability.
Error Category Percentage (%)
Limited reasoning ability of LLM 48
Incorrect response from VQA model 25
Impossible to answer by querying 11
template-based questions only
Other errors 15
@@ 6 Conclusion
In this study, we proposed a novel CoQAH methodology that answers human-written questions via a reasoning process based on a dialogue. Throughout the extensive experiments, we have shown that CoQAH achieved state-of-the-art performance on various datasets of different images (3D-rendered and chest X-ray) without finetuning using the human-written questions. We believe that this new method potentially promotes the wide applications of AI for visual question answering, improving the practicality of AI in real environment.
@@ 7 Limitation
First, we could only test a limited number of general VLM models (LLaVA [14] and GPT-4-Vision [1]) for CLEVR-Human since, currently, many LLMs do not support taking an image as an input. Second, we reported the accuracy of the AI models we ran on CLEVR-Human validation data because the test set was unavailable. Third, even though we tried our best to optimize the task instructions for the general VLM and medical foundation models, those might not be the best due to the lack of methods to tune the instructions. Fourth, we only validated CoQAH using two types of images (3D-rendered and chest X-ray) because of a limited number of template-based datasets.
@@
Supplementary Table 1: The performance of VQA models on the CLEVR testing dataset. Regardless of the models, the accuracy values were very high ($>$ 97%), meaning that those VQA models can precisely answer the template-based questions. We used the MDETR model as a template-based VQA model for CoQAH.
@@
FiLM [17] 97.7
MAC [6] 98.9
MDETR [9] 99.7
Supplementary Table 2: The performance of VQA models on the MIMIC-Diff-VQA testing dataset. Since only one VQA model (EKAID), trained with MIMIC-Diff-VQA, was available, we trained an OFA model by ourselves using a model weight of OFAbase with 182M parameters (i.e., OFA-MIMIC). Compared to the EKAID, OFA-MIMIC showed higher performance in all the metrics. We used the OFA-MIMIC as a template-based VQA model for CoQAH.
@@
EKAID [5] 0.624 0.541 0.477 0.422 0.337 0.645 1.893
OFA-MIMIC [21] 0.662 0.588 0.528 0.473 0.362 0.716 2.332
Supplementary Figure 1: Template format for the chest X-ray VQA dataset, MIMIC-Diff-VQA. A question is only composed of two parts, $\langle type\rangle$ and $\langle abnormality\rangle$. Similar to the case of the CLEVR dataset, we asked an LLM to select an option available for each $\langle type\rangle$ and $\langle abnormality\rangle$ to generate a template-based question.
Supplementary Figure 2: Task instruction of CoQAH for CLEVR.
Supplementary Figure 3: Task instruction of CoQAH for MIMIC-Diff-VQA.
Supplementary Figure 4: Task instruction of LLaVA for CLEVR-Human. (a) In the first stage, we asked LLaVA to produce an answer as a single word or phrase. (b) In some cases, LLaVA provided answers of the same meaning but different expressions. Therefore, in the second stage, we corrected them by asking LLaVA to choose one of the candidates (e.g., round for the first answer vs. sphere for the corrected answer). Note that when we gave all options for the answer at once, the performance of LLaVA was primarily degraded.
Supplementary Figure 5: Task instruction of GPT-4-Vision for CLEVR-Human. We provided a set of answers.
Supplementary Figure 6: Task instruction of Med-Flamingo for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
Supplementary Figure 7: Task instruction of MedVInT-TD for VQA-RAD and SLAKE. We limited the maximum number of output words to one for the closed-form questions and five for the open-form questions.
Supplementary Table 3: Table summarizing the answers of AI we substituted when calculating accuracy. We changed them because those words had the same meaning but different expressions, resulting in misleading performance metrics.
@@
Supplementary Figure 8: Example dialogues of CoQAH for CLEVR-Human. (a)-(c) The cases show that CoQAH correctly answered the questions and explained the reasons. (d) The answer was wrong because the template-based VQA model failed to give correct information. (e) LLM failed to query key questions and made a hasty conclusion. (f) LLM was impossible to answer the question by querying template-based questions only.
Supplementary Figure 9: Example dialogues of CoQAH for VQA-RAD and SLAKE. (a)-(c) CoQAH successfully answered the given user questions. (d) The VQA model failed to give the correct answer, and CoQAH failed accordingly. (e) The case shows that the large language model logically derived the answer even though the answer was incorrect. (f) Since the VQA model could not detect the lesion, CoQAH also gave the wrong answer.